Find the area bounded by the curve x²=y+2 and the line x+y=0.


1
Expert's answer
2021-04-26T02:03:52-0400

Let us find the x points of intersection of two graphs "y_1(x) = x^2 - 2" and "y_2(x) = - x" by solving corresponding quadratic equation "x^2 - 2 = -x". It has roots "x = -2 , x = 1".

Hence, the area between two curves is:

"S = \int_{-2}^1 (-x - (x^2 - 2 )) d x = (-\frac{x^2}{2} - \frac{x^3}{3} + 2 x)|_{-2}^1 = \frac{9}{2}".